Jack swims 9lengths of the pool in 153.9s (2 min and 33.9s), whereas Jill, the faster swimmer, covers 10lengths in the same time interval. Find the average velocity and average speed of each swimmer.
1
Expert's answer
2020-01-28T13:53:56-0500
"l=25\ m"

For Jack:

The average speed:


"v_1=\frac{9l}{t}=\frac{9(25)}{153.9}=1.46\frac{m}{s}"

The average velocity:

"u_1=\frac{l}{t}=\frac{(25)}{153.9}=0.16\frac{m}{s}"

For Jill:

The average speed:


"v_2=\frac{10l}{t}=\frac{10(25)}{153.9}=1.62\frac{m}{s}"

The average velocity:


"u_2=0\frac{m}{s}"